OpenStudy (andertj21):

What is the equation of the line that passes through (-3, -1) and has a slope of ? Put your answer in slope-intercept form. y = x + y = x - y = -x -

OpenStudy (andertj21):

Slope of 2/5

OpenStudy (anonymous):

Okay, slope intercept form is y = mx + b

OpenStudy (anonymous):

The slope is 2/5, which means we go up two, over five.

OpenStudy (andertj21):

Y=2/5x+1/5

OpenStudy (anonymous):

Sheesh, they made this one hard. If your y intercept is -3.8, what is your equation?

OpenStudy (anonymous):

-3.8 is -19/5

OpenStudy (andertj21):

Y=2/5x+1/5 Y=2/5x-1/5 This is the Answer choices

OpenStudy (anonymous):

Oh, I misread your question.

OpenStudy (anonymous):

You were right. :D

OpenStudy (andertj21):

Ok thanks:)

OpenStudy (anonymous):

No problem